OpenStudy (anonymous):

use the quadratic formula to solve the equation. If necessary, round to the nearest hundredth. x^2 + 3 = –4x (1 point) 1, 3 –1, –3 1, –3 1, –3

OpenStudy (anonymous):

x2+4x+3=0 x2+3x+x+3=0 x(x+3)+1(x+3)=0 (x+1)(x+3)=0 x= -1 or -3